OpenStudy (anonymous):

Shalika bought a purse for $120. The tax rate is 9%. Which proportion can be used to find the amount of tax Shalikia will pay?

OpenStudy (phi):

proportion: x/120 = 9/100

OpenStudy (adamk):

\[\frac{ x }{ 120 } \frac{ 9 }{ 100 }\]